Calculate no. of moles and atoms in 93 grams of phosphorus (P4)
 

Grade:9

1 Answers

Arun
5 years ago

simple first remind relation that is  mole = mass = avagrado no.

here we are taking relation b/w mole and mass then mole = mass

31 u of phosphorous combine with 1 mole = 1

1 u of phos. combine with 1 mole = 1/31

93 g of phos. combine with 1 mole = (1 / 31)ˣ93g

                                                               =  93/31

                                                                = 3

no. of mole is 3